Ad Venture is a company that produces advertising video clips. Their supply curve is Qs=-2W+6P, where
W (in 10s) is the hourly wage that the company pays their workers and P (in 1000s) is the ongoing market
price for a 1-minute commercial.
a) If the wage is $35, how much should the market price be so that Ad Venture produces 8 videos?
b) At the price you have determined at point a., how does the number of videos supplied change if the
wage decreases to $30? Show this situation in your graph.
1
Expert's answer
2019-10-22T09:18:24-0400

a) Market price for 8 videos= $2500

explanation


"Qs=-2w+6p"

"w=35\/10=3.5"


hence


"Qs=-2(3.5)+6p\n\n=-7+6p"

"p=2.5"

price="2.5*1000" "=2500"

b) 9 videos

explanation


wages $30


"Qs=-2(3)+6(2.5)=9"